OpenStudy (anonymous):

D/Dx ; Log sqrt(4-3z^2/Z+1)

OpenStudy (anonymous):

\[\log(\sqrt{\frac{4-3z}{z+1}})\] \[=\frac{1}{2}\left(\log(4-3z)-\log(z-1)\right)\] should make it easier

OpenStudy (anonymous):

I keep getting the wrong answer :/ @satellite73

OpenStudy (anonymous):

and its 3z^2

OpenStudy (anonymous):

\[=\frac{1}{2}\left(\log(4-3z^2)-\log(z-1)\right)\]

OpenStudy (anonymous):

is the log base \(e\) ?

OpenStudy (anonymous):

get \[\frac{1}{2}\left(\frac{-6z}{4-3z^2}-\frac{1}{z-1}\right)\]

OpenStudy (anonymous):

log base 5

OpenStudy (anonymous):

lol I only put -6 not -6x :/ THANK YOU